Session of pointing monitoring the 20-22 of July 2007
We check the evolution of the pointing sources used for the bolometer observations.
We observed the pointing sources doing a pointing and an Onf-Off measurements in most of the cases.
- 81 pointing sources observed.
- 72 sources have a new measured flux.
- 4 sources are found to be too faint: J0750+827, J1056+702, J1909+486, J0121+044